OpenStudy (anonymous):

how to solve log↓243 n = 4/5

OpenStudy (anonymous):

do you mean\[\log_{243} n = 4/5\] ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

think of the log function it's just means : \[243^{\frac{4}{5}} =n\]

OpenStudy (anonymous):

thanks, i forgot that step

OpenStudy (anonymous):

I'm still lost

OpenStudy (anonymous):

well \[243^{4/5}=81\]

OpenStudy (anonymous):

how do we know n=81?

OpenStudy (anonymous):

it's like saying \[\frac{243^4}{243^5}\]

OpenStudy (anonymous):

we don't I just found it :)

OpenStudy (anonymous):

hmm whoops no no it's not \[\frac{243^4}{243^5}\]

OpenStudy (anonymous):

I'm so confused

OpenStudy (anonymous):

well, you need to find n. if \[\log_{a} n= c\] then \[a^c=n\] so we have \[243^{4/5}=n\]

OpenStudy (anonymous):

yes i understand that part

OpenStudy (anonymous):

ok then \[243=3^5\] so \[243^{\frac{4}{5}}=3^4=81\]

OpenStudy (anonymous):

so you simplified 245 and got 3^4, so the fraction 4/5 is forgotten, and you just simply make 81 your answer?

OpenStudy (anonymous):

yep!

OpenStudy (anonymous):

lol thanks :)

OpenStudy (anonymous):

n=81

OpenStudy (anonymous):

thats what my answer key says also

OpenStudy (anonymous):

seems likely :)